OpenStudy (anonymous):

3. Solve, using the substitution method. 5x + 2y = 7 y = x – 7

TheSmartOne (thesmartone):

Just plug in y=x-7 into 5x+2y=7

TheSmartOne (thesmartone):

so 5x+2(x-7)=7 now solve for x...

TheSmartOne (thesmartone):

You would distribute the 2 inside the paranthesis and add like terms and find x then you would plug in x and solve for y...

TheSmartOne (thesmartone):

@helpdummy

OpenStudy (anonymous):

ok so then how would i solve for y when all there is is x's @TheSmartOne

TheSmartOne (thesmartone):

You would find y by plugging back what you get from x.

OpenStudy (anonymous):

@TheSmartOne i am thinking its no solution

TheSmartOne (thesmartone):

5x+2x-14=7 7x-14=7 7x=21 x=3

TheSmartOne (thesmartone):

Now plug back in x into any equation y=3-7 y=-4

TheSmartOne (thesmartone):

So your solution is x=3 and y=-4

OpenStudy (anonymous):

you didnt have to answer because i feel bad i didnt do the work myself

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

its okay..